OpenStudy (rational):

does this question make sense Phillip received 75 points on a project for school. He can make changes and receive two-tenths of the missing points back. He can do this 10 times. Create the formula for the sum of this geometric series, and explain your steps in solving for the maximum grade Phillip can receive.

zepdrix (zepdrix):

What is the amount of the missing points? Are we assuming it's a 100 point project?

OpenStudy (rational):

not given i think that's one of the careless mistakes from teacher, but even if i assume it is 100 points project, im not able to work the solution...

zepdrix (zepdrix):

And then I guess each 2/10 is being applied to the "new missing points" .. or something? strange problem 0_o

OpenStudy (rational):

yeah exactly thats what troubled me if they mean the net missing points, then Philip can achieve 99+ easily in 10 submissions i think..

zepdrix (zepdrix):

hmm yah it seems to be a poorly written problem D: kinda confusing. I guess if you make a buuuunch of assumptions you can make the problem work out ok...\[\Large\rm 75+\left(\frac{2}{10}\right)m+\left(\frac{2}{10}\right)^2m+...+\left(\frac{2}{10}\right)^{10}m\]Where m is the missing amount (i'm assuming it's 25 to start). and then ya do some kind of summation thing :d weirrrrd

OpenStudy (rational):

Nice :) I think thats what the teacher wants! I answered the same here but yeah all this is not agreeing with the question as it stands... http://gyazo.com/422e993848b64985157e22ca4e8f9531

zepdrix (zepdrix):

Oh I screwed something up there didn't I.. If you apply (2/10)m, then the remaining missing amount is (8/10)m So wouldn't the next piece be (2/10)(8/10)m??

OpenStudy (rational):

Yeah would the series still remain geometric if we change the factor 2/10 to 8/10

zepdrix (zepdrix):

the 6.25 does LOOK LIKE what the teacher was probably looking* for.. but boy the problem doesn't seem to line up with that :O

OpenStudy (amistre64):

25*.8 is the remaining points after a submission.

OpenStudy (amistre64):

missing = m = 25(.8) reward = m(.2)

OpenStudy (amistre64):

for submission points, the set of terms p p1 = 25(.8)^0 (.2) p2 = 25(.8)^1 (.2) p3 = 25(.8)^2 (.2) right?